Circular No. 7699

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001do IN UGC 11459 R. Chornock, M. Modjaz, and A. V. Filippenko, University of Cal-

ifornia at Berkeley (UCB), report that inspection of CCD spectra (range 330–1000 nm), obtained on Aug. 23 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2001do (IAUC 7682) is an unusual supernova of type II. The spectrum exhibits low-contrast absorption and emission features of H, Fe, and Na superimposed on a smooth continuum with a color temperature of ∼ 5800 K.

SUPERNOVAE 2001dk, 2001dq, 2001ds Chornock et al. report that inspection of CCD spectra, obtained as

above, reveals that SN 2001dk (IAUC 7675, 7676) is a normal type II-P supernova, a few weeks past maximum light. Well-developed P-Cyg lines of H, Fe, Sc, and Ca are present. The expansion velocity measured from the Hα absorption is 9000 km/s. Additional spectra show that SN 2001ds (IAUC 7684) and SN 2001dq (IAUC 7683) are supernovae, probably of type Ia (∼ 3 weeks past maximum brightness) and type Ic, respectively.

V1178 SCORPII D. K. Lynch and R. J. Rudy, The Aerospace Corporation; and J. C.

Wilson, Cornell University, report 0.8–2.5-μm spectrophotometry of V1178 Sco (= N Sco 2001) using the Palomar 1.5-m telescope (+ CorMASS nearinfrared spectrograph) on July 30.16 UT, ∼ 40 days after peak brightness (June 20): “The object has formed large amounts of dust in the 19 days between observations (see IAUC 7679). It has faded by over 4 magnitudes in the J band (1.25 μm), and its spectrum shows evidence of both heavy reddening and significant thermal emission beyond 2 μm. Based on our spectra and the light curve, it appears as though dust formation may have begun around July 10, ∼ 3 weeks after peak brightness. Magnitudes derived from the spectrophotometry: July 11, J = 9.2, H [1.65 μm] = 8.11, K(short) [2.16 μm] = 6.8; July 30, 13.3, 10.6, 8.4.”

SUPERNOVA 2001dz IN UGC 471 Corrigendum. On IAUC 7696, lines 6–7, the offset should read 13′′ east

and 7′′ south of the center of UGC 471.

2001 August 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green